OpenStudy (anonymous):

Suppose that the x_intercepts of the graph of y = f(x) are -3 and 2. Then the x-intercepts of the graph of the graph of y = 2f(x) are -3 and 2. true or false

OpenStudy (zzr0ck3r):

notice that \(f(-3)=f(2)=0\) implies that \(2f(-3)=2f(2)=2*0=0\)

OpenStudy (zzr0ck3r):

So \(2f(-3)=0\) and \(2f(2)=0\)

OpenStudy (zzr0ck3r):

So the roots of \(2f(x)\) are ?